# latexOverwriting file

### v1

@@
To Appear In SIAM News
[…]
(February 2, 2024)
@@ 1 Introduction
Estimating the statistics of the state of a dynamical system, from partial and noisy observations, is both mathematically challenging and finds wide application. Furthermore, the applications are of great societal importance, including problems such as probabilistic weather forecasting [12] and prediction of epidemics [13]. Particle filters provide a well-founded approach to the problem, leading to provably accurate approximations of the statistics [6]. However these methods perform poorly in high dimensions [1, 20]. In 1994 the idea of ensemble Kalman filtering was introduced [8] leading to a methodology that has been widely adopted in the geophysical sciences [21] and also finds application to quite general inverse problems [10]. However, ensemble Kalman filters have defied rigorous analysis of their statistical accuracy, except in the linear Gaussian setting [15, 16]. In this article we describe recent work which takes first steps to analyze the statistical accuracy of ensemble Kalman filters beyond the linear Gaussian setting [3]. The subject is inherently technical, as it involves the evolution of probability measures according to a nonlinear and nonautonomous dynamical system; and the approximation of this evolution. It can nonetheless be presented in a fairly accessible fashion, understandable with basic knowledge of dynamical systems, numerical analysis and probability.
@@ 2 Filtering Problem
$v_{0}\perp\!\!\!\!\perp\{\xi_{n}\}_{n\in\mathbb{Z}^{+}}\perp\!\!\!\!\perp\{\eta_{n+1}\}_{n\in\mathbb{Z}^{+}}.$
The objective of filtering is to determine the probability distribution on $v_{n}$ given all the data acquired up to that point. To this end we define, for a given realization of the data, denoted by a dagger ${}^{\dagger}$,
$Y^{\dagger}_{n}=\{{y}^{\dagger}_{\ell}\}_{\ell=1}^{n},\qquad v_{n}|Y^{\dagger}_{n}\sim{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{n}.$
With this notation we can state the objective of filtering more precisely: it is to determine the filtering distribution or true filter ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{n}$ and update it sequentially in $n$. As we now show ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{n}$ evolves according to a nonautonomous and nonlinear dynamical system on the space of probability measures.
@@
For simplicity we will use the same symbol for measures and their densities throughout this article. Before writing the dynamical system for ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{n}$ we first determine the evolution of the measures ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\pi}}_{n},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{r}}_{n}$ defined by
$v_{n}$ $\sim{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\pi}}_{n},\qquad(v_{n},y_{n})\sim{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{r}}_{n}.$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\pi}}_{n+1}$ $=\mathsf{P}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\pi}}_{n},$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{r}}_{n+1}$ $=\mathsf{Q}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\pi}}_{n+1}$
@@
$\mathsf{P}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\pi}}(u)=\frac{1}{\sqrt{(2\pi)^{d}\det\Sigma}}\int\exp\left(-\frac{1}{2}|u-\Psi(v)|_{\Sigma}^{2}\right){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\pi}}(v)\,\mathrm{d}v.$
@@
$\mathsf{Q}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\pi}}(u,y)=\frac{1}{\sqrt{(2\pi)^{K}\det\Gamma}}\exp\left(-\frac{1}{2}\bigl{\lvert}y-h(u)\bigr{\rvert}_{\Gamma}^{2}\right){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\pi}}(u).$
@@
The evolution for ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\pi}}_{n}$, which describes the probability of the state $v_{n}$, is determined by a Markov process on $\mathbb{R}^{d}$ defined via the linear operator $\mathsf{P}.$ The linear operator $\mathsf{Q}$ lifts ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\pi}}_{n}$ to the joint space of state and data $(v_{n},y_{n}).$ It is worth highlighting that, by moving from the evolution of $(v_{n},y_{n})\in\mathbb{R}^{d}\times\mathbb{R}^{K}$ (finite dimensions) to the evolution of $({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\pi}}_{n},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{r}}_{n})\in\mathcal{P}(\mathbb{R}^{d}\times\mathbb{R}^{K})$ (infinite dimensions) we have converted a nonlinear stochastic problem into a linear autonomous one.
@@
However, the dynamical system for ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{n}$ is nonlinear and nonautonomous. To determine the dynamical system for ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{n}$ we need to introduce a nonlinear operator on probability measures. Specifically, let $\mathsf{B}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};{y}^{\dagger})\colon\mathcal{P}(\mathbb{R}^{d}\times\mathbb{R}^{K})\to\mathcal{P}(\mathbb{R}^{d})$ describe conditioning of a joint random variable $(v,y)$ on observation $y={y}^{\dagger}$:
@@
$\mathsf{B}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}};{y}^{\dagger})(u)=\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}}(u,{y}^{\dagger})}{\int_{\mathbb{R}^{d}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}}(u,{y}^{\dagger})\,\mathrm{d}u}.$
@@
Armed with this we define $v_{n+1}|Y^{\dagger}_{n}\sim\widehat{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}}_{n+1}$ and observe that
@@
$\widehat{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}}_{n+1}$ $=\mathsf{P}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{n},\qquad$ $v_{n+1}|Y^{\dagger}_{n}\sim\widehat{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}}_{n+1}$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}}_{n+1}$ $=\mathsf{Q}\widehat{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}}_{n+1},\qquad$ $(v_{n+1},y_{n+1})|Y^{\dagger}_{n}\sim{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}}_{n+1}$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{n+1}$ $=\mathsf{B}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\rho}}_{n+1};{y}^{\dagger}_{n+1}),\qquad$ ${\rm conditioning}\,{\rm on}\,y_{n+1}={y}^{\dagger}_{n+1}.$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{n+1}$ $=\mathsf{B}(\mathsf{Q}\mathsf{P}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{n};{y}^{\dagger}_{n+1}),\qquad\mu_{0}=\mathcal{N}(m_{0},C_{0}).$
This evolution may be thought of in terms of sequential application of Bayes theorem: $\mathsf{P}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{n}$ is prior prediction; $\mathsf{L}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};{y}^{\dagger}):=\mathsf{B}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};{y}^{\dagger})\circ\mathsf{Q}$ maps prior to posterior according to Bayes theorem. This leads to the following equivalent formulation of the evolution:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{n+1}$ $=\mathsf{L}(\mathsf{P}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{n};{y}^{\dagger}_{n+1}),\qquad\mu_{0}=\mathcal{N}(m_{0},C_{0}).$
The resulting evolution is an infinite dimensional problem since ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@rgb@stroke{0}{0}{0}\pgfsys@color@rgb@fill{0}{0}{0}\mathfrak{\mu}}_{n}$ is a probability density over $\mathbb{R}^{d}.$ Approximating this evolution is thus a significant challenge.

### v3

# Statistical Accuracy of Approximate Filtering Methods

@@
J. A. Carrilloa, Mathematical Institute, University of Oxford, UK F. Hoffmannb, Department of Computing and Mathematical Sciences, Caltech, USA A. M. Stuart c, Department of Computing and Mathematical Sciences, Caltech, USA U. Vaes d, MATHERIALS project-team, Inria Paris, France CERMICS, École des Ponts, France
(May 31, 2025)
@@ 1 Introduction
Estimating the statistics of the state of a dynamical system, from partial and noisy observations, is both mathematically challenging and finds wide application. Furthermore, the applications are of great societal importance, including problems such as probabilistic weather forecasting [kalnay2003atmospheric] and prediction of epidemics [keeling2005networks]. Particle filters provide a well-founded approach to the problem, leading to provably accurate approximations of the statistics [doucet2001sequential]. However these methods perform poorly in high dimensions [MR2459233, snyder2008obstacles]. In 1994 the idea of ensemble Kalman filtering was introduced [evensen1994sequential] leading to a methodology that has been widely adopted in the geophysical sciences [van2019particle] and also finds application to quite general inverse problems [iglesias2013ensemble]. However, ensemble Kalman filters have defied rigorous analysis of their statistical accuracy, except in the linear Gaussian setting [le2009large, mandel2011convergence]. In this article we describe recent work which takes first steps to analyze the statistical accuracy of ensemble Kalman filters beyond the linear Gaussian setting [carrillo2022ensemble]. The subject is inherently technical, as it involves the evolution of probability measures according to a nonlinear and nonautonomous dynamical system; and the approximation of this evolution. It can nonetheless be presented in a fairly accessible fashion, understandable with basic knowledge of dynamical systems, numerical analysis and probability. We undertake such a presentation here.
@@ 2 Filtering Problem
$v_{0}\perp\!\!\!\!\perp\{\xi_{n}\}_{n\in\mathbb{Z}^{+}}\perp\!\!\!\!\perp\{\eta_{n+1}\}_{n\in\mathbb{Z}^{+}}.$ (2.1)
The objective of filtering is to determine the probability distribution on $v_{n}$ given all the data acquired up to that point. To this end we define, for a given realization of the data, denoted by a dagger ,
$Y^{\dagger}_{n}=\{{y}^{\dagger}_{\ell}\}_{\ell=1}^{n},\qquad v_{n}|Y^{\dagger}_{n}\sim{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{n}.$
With this notation we can state the objective of filtering more precisely: it is to determine the filtering distribution or true filter ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{n}$ and update it sequentially in $n$. As we now show ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{n}$ evolves according to a nonautonomous and nonlinear dynamical system on the space of probability measures.
@@
For simplicity we will use the same symbol for measures and their densities throughout this article. Before writing the dynamical system for ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{n}$ we first determine the evolution of the measures ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\pi}}_{n},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{r}}_{n}$ defined by
$v_{n}$ $\sim{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\pi}}_{n},\qquad(v_{n},y_{n})\sim{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{r}}_{n}.$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\pi}}_{n+1}$ $=\mathsf{P}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\pi}}_{n},$
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{r}}_{n+1}$ $=\mathsf{Q}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\pi}}_{n+1}$
@@
$\mathsf{P}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\pi}}(u)=\frac{1}{\sqrt{(2\pi)^{d}\det\Sigma}}\int\exp\left(-\frac{1}{2}|u-\Psi(v)|_{\Sigma}^{2}\right){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\pi}}(v)\,\mathrm{d}v.$
@@
$\mathsf{Q}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\pi}}(u,y)=\frac{1}{\sqrt{(2\pi)^{K}\det\Gamma}}\exp\left(-\frac{1}{2}\bigl{\lvert}y-h(u)\bigr{\rvert}_{\Gamma}^{2}\right){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\pi}}(u).$
@@
The evolution for ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\pi}}_{n}$, which describes the probability of the state $v_{n}$, is determined by a Markov process on $\mathbb{R}^{d}$ defined via the linear operator $\mathsf{P}.$ The linear operator $\mathsf{Q}$ lifts ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\pi}}_{n}$ to the joint space of state and data $(v_{n},y_{n}).$ It is worth highlighting that, by moving from the evolution of $(v_{n},y_{n})\in\mathbb{R}^{d}\times\mathbb{R}^{K}$ (finite dimensions) to the evolution of $({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\pi}}_{n},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{r}}_{n})\in\mathcal{P}(\mathbb{R}^{d}\times\mathbb{R}^{K})$ (infinite dimensions) we have converted a nonlinear stochastic problem into a linear autonomous one.
@@
However, the dynamical system for ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{n}$ is nonlinear and nonautonomous. To determine the dynamical system for ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{n}$ we need to introduce a nonlinear operator on probability measures. Specifically, let $\mathsf{B}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};{y}^{\dagger})\colon\mathcal{P}(\mathbb{R}^{d}\times\mathbb{R}^{K})\to\mathcal{P}(\mathbb{R}^{d})$ describe conditioning of a joint random variable $(v,y)$ on observation $y={y}^{\dagger}$:
@@
$\mathsf{B}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}};{y}^{\dagger})(u)=\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}}(u,{y}^{\dagger})}{\int_{\mathbb{R}^{d}}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}}(u,{y}^{\dagger})\,\mathrm{d}u}.$
@@
Armed with this we define $v_{n+1}|Y^{\dagger}_{n}\sim\widehat{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}}_{n+1}$ and observe that
@@
$\widehat{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}}_{n+1}$ $=\mathsf{P}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{n},\qquad$ $v_{n+1}|Y^{\dagger}_{n}\sim\widehat{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}}_{n+1}$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}}_{n+1}$ $=\mathsf{Q}\widehat{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}}_{n+1},\qquad$ $(v_{n+1},y_{n+1})|Y^{\dagger}_{n}\sim{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}}_{n+1}$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{n+1}$ $=\mathsf{B}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\rho}}_{n+1};{y}^{\dagger}_{n+1}),\qquad$ ${\rm conditioning}\,{\rm on}\,y_{n+1}={y}^{\dagger}_{n+1}.$
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{n+1}$ $=\mathsf{B}(\mathsf{Q}\mathsf{P}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{n};{y}^{\dagger}_{n+1}),\qquad\mu_{0}=\mathcal{N}(m_{0},C_{0}).$
This evolution may be thought of in terms of sequential application of Bayes theorem: $\mathsf{P}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{n}$ is prior prediction; $\mathsf{L}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};{y}^{\dagger}):=\mathsf{B}(\mathord{\color[rgb]{0.67,0.67,0.67}\definecolor[named]{pgfstrokecolor}{rgb}{0.67,0.67,0.67}\pgfsys@color@gray@stroke{0.67}\pgfsys@color@gray@fill{0.67}\bullet};{y}^{\dagger})\circ\mathsf{Q}$ maps prior to posterior according to Bayes theorem. This leads to the following equivalent formulation of the evolution:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{n+1}$ $=\mathsf{L}(\mathsf{P}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{n};{y}^{\dagger}_{n+1}),\qquad\mu_{0}=\mathcal{N}(m_{0},C_{0}).$
The resulting evolution is an infinite dimensional problem since ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\mathfrak{\mu}}_{n}$ is a probability density over $\mathbb{R}^{d}.$ Approximating this evolution is thus a significant challenge.